CAN SOMEONE PLEASE HELP ME WITH THIS PROBLEM:
A triangle has a perimeter of 96 inches. Its shortest side measures 3 inches shorter than its middle side, and longest side measures 12 inches longer than its middle side. Find the lengths of the triangle's three side.
: CAN SOMEONE PLEASE HELP ME WITH THIS PROBLEM:
A triangle has a perimeter of 96 inches. Its shortest side measures 3 inches shorter than its middle side, and longest side measures 12 inches longer than its middle side. Find the lengths of the triangle's three side.

Answer by checkley77(3848)

S=M-3
L=M+12
S+M+L=96
(M-3)+M+(M+12)=96
3M+9=96
3M=96-9
3M=87
M=87/3
M=29 ANSWER FOR THE MIDDLE SIDE.
S=29-3=26 FOR THE SHORT SIDE.
L=29+12=41 FOR THE LONG SIDE.
PROOF:
29+26+41=96
96=96